IceClaws3309:

Rewrite the expression with rational exponents as a radical expression by extending the properties of integer exponents. (64x^2/3)^-1/3 the square root of three to the one sixth power the ninth root of three squared the square root of three to the sixth power the sixth root of three

jhonyy9:

use this \[\left[ \left( 64x^{2/3} \right)^{1/3} \right]^{-1} = ?\]

jhonyy9:

what you need to know again that the denominator of fractional exponent is the index of radical so what mean for example \[x^{2/3} = \sqrt[3]{x^{2}}\]

IceClaws3309:

tysm